OpenStudy (anonymous):

Refer to the definition of a dilution ratio as explained below to determine the ratio. A dilution ratio as used in a laboratory is a very specific ratio. The numerator is the amount of substance being diluted and the denominator is the total amount of the mixture, including the diluting agent and the substance being diluted. Calculate the dilution ratio of a solution of 75 milliliters of chlorine bleach diluted with 465 milliliters of distilled water. Write the ratio in simplest form using a fraction.

jimthompson5910 (jim_thompson5910):

75 milliliters of chlorine bleach 465 milliliters of distilled water 465+75 = 540 mL of mixture (chlorine bleach + water)

jimthompson5910 (jim_thompson5910):

we then divide the amount of pure chlorine bleach (75) by the amount of mixture (540) 75/540 = 5/36 don't forget to reduce fully

OpenStudy (anonymous):

5/36 is reduced

jimthompson5910 (jim_thompson5910):

btw, 5/36 = 0.1388888 roughly, so it's 13.88% bleach

jimthompson5910 (jim_thompson5910):

yes I meant when going from 75/540 to 5/36

OpenStudy (anonymous):

so i just write 5/36

jimthompson5910 (jim_thompson5910):

well you show all the work too

OpenStudy (anonymous):

its just 5/36

jimthompson5910 (jim_thompson5910):

5/36 is the final answer, but you should show steps

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

np